Find the slope of the line passing through the given pair of points.(8, -6) and (3, -6)
Answer by stanbon(75887)

Find the slope of the line passing through the given pair of points.(8, -6) and (3, -6)
slope = (-6--6)/(8-3) = 0/5 = 0
-----------------------------------
Cheers,
Stan H.
===========